# Mixture of Experts in Image Classification: What’s the Sweet Spot?

### v1

@@ Introduction
Architecture MoE ×$\times 10^{6}$ Sample FLOPs IN-1K
@@
Architecture # experts #Params Per samples IN-1K
(×$\times 10^{6}$) #Paramsact Top 1 acc.
ConvNext-T no MoE 28.6 28.6 82.1
4 54.3 25.5 82.1
8 47.4 25.5 82.1
ConvNext-T 16 46.3 25.5 81.7
ViT-S no MoE 22.0 22.0 79.9
4 29.1 25.0 79.8
8 38.6 25.0 80.5
ViT-S 16 57.5 25.0 80.2
@@
MoE strategy #Params Per samples FLOPs IN-1K
(×$\times 10^{6}$) #Paramsact Top 1 acc.
@@
Table 6: ImageNet-21k training on ConvNext-T: comparison between various numbers of experts, with a MLP ratio of 4.
Table 3 displays the effects of various placements of MoE layers, comparing the following configurations: (a) Every 2: a MoE layer replaces every second layer; (b) Stage: a MoE layer replaces the final layer of each stage, resulting in four MoE layers throughout the network for ConvNext; (c) Last 2: a MoE layer replaces the final layer of each of the last two stages; (d) Last 3: Same as “Last 2”, but with an additional MoE layer inserted in the middle of stage 3.
As shown in Table 3, “Last 2” strategy is the most robust choice: it performs well across all architectures. On the opposite, “Every 2” performs worst for ConvNext architecture, and best for ViT. Consequently, in the following, all experiments use “Last 2” for ConvNext and “Every 2” for ViT.
Influence of the number of experts
Table 4 presents the performance of ConvNext and ViT architectures on ImageNet-1K training with the “Last 2” strategy MoE layers. It provides a comprehensive view of how the number of experts influences the size (parameter count) of the models, and their Top-1 accuracy on this dataset.
The experimental results suggest that four experts yield the best performance for ConvNext, while eight experts are optimal for ViT. This is showcased by the Top-1 accuracy rates, which are consistent at 82.1% for ConvNext with four experts, and reach a peak of 80.5% for ViT with eight experts. However, we note that increasing the number of experts to 16 has a detrimental effect on both architectures. Specifically, for ConvNext, the top-1 accuracy slightly drops to 81.7%, and for ViT, the performance plateaus at 80.2%.
These findings highlight that while MoE layers can enhance model performance, there is a delicate trade-off to be struck in terms of the number of experts. Exceeding the optimal number can lead to suboptimal results, negating the potential benefits of the MoE integration. Furthermore, in our detailed analysis, we explore the interplay between the number and the size of each expert. Our investigation revealed that, for isotropic networks such as ViT and ConvNext iso., reducing the size of experts adversely affects performance, while for ConvNext, it has no significant impact. Moreover, during the course of our experiments, we observed that while the Top-1 configuration was superior for ConvNext, the Top-2, for ViT configuration was comparable or even better than the ConvNext Top-1 configuration, leading us to employ Top-1 for ConvNext and Top-2 for ViT.
Results on ImageNet-21k
MoE strategy #Params Per samples FLOPs IN-1K
(×$\times 10^{6}$) #Paramsact Top 1 acc.
no MoE 28.6 28.6 4.5G 82.9
every 2 97.5 28.6 4.5G 82.9
stage 78.2 28.6 4.5G 83.5
last 3 70.0 28.6 4.5G 83.5
last 2 70.0 28.6 4.5G 83.5
last 2 mlp_ratio 2 46.3 25.6 4.2G 83.4
# experts #Params Per samples FLOPs IN-1K
(×$\times 10^{6}$) #Paramsact Top 1 acc.
no MoE 28.6 28.6 4.5G 82.9
8 54.3 28.6 4.5G 83.5
16 117.1 28.6 4.5G 83.6
32 211.6 28.6 4.5G 83.4
Table 8: ImageNet-21k training on ConvNext-T: comparison between various numbers of experts, with a MLP ratio of 4.
Tables 8 and 8 present the results of models trained on ImageNet-21k, reporting ImageNet-1k accuracy after fine-tuning We analyse the effect of MoE layer positioning, expert size (Table 8), and expert count (Table 8). The results demonstrate that, when large volumes of data are accessible, a greater number of experts can be effectively deployed. Notably, the use of sixteen experts, which was sub-optimal for ImageNet-1K, does not negatively affect performance on ImageNet-21k. This suggests that increasing the number of experts together with the volume of data could lead to valuable enhancements.
In terms of MoE layer configuration, our data showed no performance improvements with the addition of more layers, and the “every 2” strategy again yielded the poorest results for ConvNext based architecture.
Discussion
Hierarchical vs isotropic models
First, one can note that the best MoE designs typically apply experts only to the last layers. ViT operates at the same resolution at each layer after the input images have been cut into patches. This isotropic design implies that MoE is applied at layers which are not particularly large. On the other hand, with hierarchical models such as ConvNext, applying MoE to the last layers considerably increases model size. Indeed, as presented in Table 8, employing eight experts doubles the size of the network in spite of using only two layers of MoE.
Positions and numbers of MoE layers
There are two prevalent strategies in the literature concerning the position of MoE layer: “Every 2” and “Last 2”. In the context of hierarchical architectures such as ConvNext, the “Every 2” strategy dramatically increases the number of weights while simultaneously yielding inferior results. However, for ViT “Last 2” and “Every 2” are two viable strategies. This is inline with the results of V-MoE (Riquelme et al. 2021) on large-scale datasets. Overall, our results confirm that “Last 2” is a solid starting point, as it consistently yields positive outcomes across all tested architectures. However, it may not always be the optimal choice for every architecture. For instance, in the case of ViT, “Every 2” demonstrates better results. Regarding the number of experts, we noted that model performance quickly saturates with the number of experts, as reproduced in Table 8 (more details on Figure 5).
Evaluating MoE in image classification
With the ImageNet benchmark increasingly reaching saturation, the focus in contemporary computer vision models has shifted towards computational efficiency and scaling prowess. An essential question arises: Does the MoE model enhance this aspect in image classification?
Our experiments with ImageNet reveal an intriguing insight. While integrating MoE into models like ConvNext and ViT does enhance performance, this improvement tends to plateau in models exceeding 100M parameters. This observation aligns with existing literature, as evidenced by results presented in Figure 2 (See Table 14 and Figure 8). One can observe that as the base dense network increases, the performance disparity between MoE and its dense counterpart decreases.
Exploring this further in large-scale datasets, such as JFT 3 billion (Zhai et al. 2022), reveals similar findings. For instance, the vision transformer model SoViT, with 400M parameters (Alabdulmohsin et al. 2023), exhibits comparable performance to the 15B V-MoE (Riquelme et al. 2021) when both use ViT-based architectures. Similarly, when considering the JFT-3B dataset (Zhai et al. 2022), the 5.6B LiMoE-H model (Mustafa et al. 2022) and Coca-Large (Yu et al. 2022) lead to analogous results when evaluated on a per-sample parameter basis.
To summarize, while the MoE model does not appear to redefine the state-of-the-art when matched against dense models on both ImageNet and billion-scale datasets, its real strength emerges in enhancing smaller models. By integrating MoE, these compact models can achieve better performance, pushing their capabilities closer to the forefront of current benchmarks. Yet, it is essential to acknowledge that, in the broader context, MoE might not drive significant advancements in the overall state-of-the-art. This nuanced perspective is further enforced in the appendix, with the data showcased in Table 14.
Figure 2: Improvement obtained thanks to MoE (y-axis) vs accuracy without MoE (x-axis), for different pretraining sizes and specifically for image classification. The more accurate the model, the lower the impact. The larger the pretraining, the greater the impact. For each pretraining size, a dashed line shows for each accuracy without MoE the maximum improvement obtained for at least this accuracy. Results are extracted from Tables 2, 1 and 14 (in Appendix).
Robustness Evaluation
We evaluated MoE models, trained or fine-tuned on ImageNet-1k, against a range of robustness benchmark datasets, such as ImageNet-A (IMA) (Hendrycks et al. 2021b), ImageNet-R (IMR) (Hendrycks et al. 2021a), ImageNet-Sketch (IMSK) (Wang et al. 2019), and ImageNet V2 (IMV2) (Recht et al. 2019), the latter being used as a measure of overfitting. Table 9 reports Top-1 accuracy for all datasets.
@@ Discussion
Table 9: Robustness of MoE w.r.t. to domain change, for two different pretrainings both fine-tuned on ImageNet-1k: Top-1 accuracies on various datasets.
@@
Table 9 shows that MoE models underperform their dense counterparts on ImageNet-1k. These results seem to contradict the ones reported by Li et al. (2022) in which MoE models appear as natural domain generalizers. We note that MoE models are better for the robustness metric, they were actually already better on ImageNet-1k. So, these results actually emphasize the success of MoE when the model is sufficiently small for being improved by a MoE, as discussed in the section “Evaluating MoE in Image Classification”: the domain generalization holds only in the regime in which MoE improves the base results.
Model Inspection
@@
The objective of this section is to scrutinize the routing component of the model. We will present the top expert for each routing layer. Note that when considering the top-2 configuration, we only report, for this visualization, the top-1 of these two experts.
Visual Inspection. Figure 3 in appendix, provides a visualization of the portions of images that are assigned to the different experts. This reveals that some experts specialize in specific elements of an image, such as animals, buildings, or objects. This trend is especially apparent when considering the top-1 MoE, but the distinction becomes less clear with the top-2 MoE. In both instances, the clusters formed by each expert are challenging to interpret, and the spatial locations of each expert tend to lack clarity and can be quite ambiguous.
Number of experts involved per image. We quantified the average number of experts contributing to processing an image. In the initial layers, the majority of experts are typically engaged for each image. With fewer than eight experts, most are utilized, even in deeper layers. For instance, in the ViT model with eight experts, an average of seven experts contributed per image, while in the ConvNext model, they are six. As we increase the number of available experts per MoE layer, the number of contributing experts per image also rises. For example, in ConvNext with 16 experts, an average of 10.5 experts are involved, and with 32 experts, this number increases to 15.8. This suggests that each expert is likely focused on a limited number of image patches. This observation is confirmed by the cumulative distribution function (CDF) of each expert per image, displayed in the appendix (see Figure 7 in the Appendix).
Correlation Between Experts and Labels. Figure 4 (top) in the Appendix presents the correlation between the number of expert occurrence and individual classes in ImageNet-1k. The class IDs in ImageNet-1k are organized in such a way that adjacent classes often share similar attributes. For instance, class IDs ranging from 151 to 268 are all dedicated to different dog breeds. From Figure 4, we note that the occurrences in the first three MoE layers do not match the ImageNet-1k classes, with most experts appearing uniformly across different classes. However, in the last three MoE layers, there is a noticeable trend of more experts aligning with specific ImageNet-1k classes. This distinction is especially pronounced between experts focusing on animals (classes up to ID 397, such as experts 0 and 6 in Figure 4-(e)) and those centred on objects (classes beyond ID 397, like experts 1 and 5 in Figure 4-(e)). As the number of experts increases, this class-specific alignment still holds, though it remains challenging to clearly define the roles of these experts since they often span several scattered classes rather than forming tight, contiguous groups that would fit each class, as illustrated in Figure 6 in the appendix.
Expert Similarity. In order to quantify the similarity between experts, we use a method derived from the Jaccard index. For each image, we tally the occurrences $c_{i}$ of each expert $E_{i}$. The similarity between two experts, $E_{i}$ and $E_{j}$, is defined as $S_{ij}=\frac{|E_{i}\cap E_{j}|}{|E_{i}\cup E_{j}|}=\frac{\sum^{E_{i}\cap E_{j}}c_{i}+c_{j}-|c_{i}-c_{j}|}{\sum^{E_{i}\cup E_{j}}c_{i}+c_{j}}$, where ${E_{i}\cap E_{j}}$ represents the images where both $E_{i}$ and $E_{j}$ appear together, while ${E_{i}\cup E_{j}}$ refers to all images where either $E_{i}$ or $E_{j}$ is present at least once. The absolute value term $|c_{i}-c_{j}|$ acts as a corrective factor to account for scenarios where one expert may be deployed for a minimal number of patches while the other is utilized for the majority, indicating a lack of synergy between the two experts. In the bottom panel of Figure 4, we observe a distinct pattern in the utilization of experts across the stages of the MoE. In the initial layers, there is a prevalent co-occurrence of multiple experts, which highlights a concentration on local image features across all the images. Conversely, as we analyse deeper into the network, the frequency of expert co-occurrence diminishes. This shift indicates a transition in focus from local image elements to more semantically rich aspects, with individual experts concentrating on specific semantic attributes unique to certain images.
Are experts well partitioned, clustered?
Firstly, for a general task such as ImageNet-1k, there is no clear, natural decomposition into experts, making it difficult to precisely assess if one expert decomposition is better than another, aside from comparing final accuracy. This comparison, however, provides limited insight into the quality of the experts. The literature (Chen et al. 2023; Mittal, Bengio, and Lajoie 2022; Mohammed, Liu, and Raffel 2022) suggests that, without explicit enforcement, MoE models struggle to separate tasks or find natural clusters beneath them. For instance, when a task separation was explicitly incorporated into the loss, there was a marked improvement in both accuracy and the ease with which experts could be pruned for specific tasks (Chen et al. 2023). Similarly, Mittal, Bengio, and Lajoie (2022) found that the enforced decomposition resulted in substantial gains, while the absence of such enforcement left the model unable to discover optimal partitions.
Consequently, we make three observations on ImageNet. (a) As the number of experts increases, there is a corresponding rise in the number of experts actively involved. Even with 32 experts, almost half of the experts are used per image in the last MoE layers. This seems counter-intuitive, since ImageNet is centred around objects, it should plateau at some point, at least for deep layers. Instead, experts are not well aligned with classes and spread widely, as depicted in Figure 4. (b) An analysis of the cumulative distribution function (CDF) of experts’ engagement (Figure 7 in the Appendix) reveals a prevalent trend where most experts are assigned to less than 20 patches of an image, equating to less than 10% of the total image area, 50% of the time. Moreover, these patches are not necessarily contiguous. (c) When we look at visualization of each expert like in Figure 3, we observe a concentration of experts in the image for MoE layers near the output. The interpretation of each expert’s role becomes challenging, and each expert does not seem to adhere closely to each object or to each part of the object location.
Conclusion
Implementing Mixture-of-Experts (MoE) models for computer vision tasks remains an open and challenging task. Our experimental findings indicate that using MoE in computer vision, particularly for large models, does not substantially surpass the current state-of-the-art in accuracy, robustness, domain generalization, or per-sample complexity. This observation applies to both convolutional networks and vision transformers. These results highlight a contrast between text processing and computer vision applications of MoE. While, in the text domain tasks, MoE models benefit from adding weighted components, enhancing knowledge storage within the model, in computer vision tasks, they predominantly hinge on extrapolation and feature extraction, where MoE models seem to offer limited advantages. Interestingly, we observe the most notable improvements in smaller-scale models, especially in image classification tasks, where the number of activated parameters per sample is low. In these cases, MoE layers provided moderate enhancements, suggesting that the effectiveness of MoE models in computer vision may be context- and scale-dependent. For instance, MoE models show potential benefits in scenarios with fewer activations per sample, in contrast with dense (non-MoE) models counterparts.
Appendix A Exploration of routing architectures
We investigated three distinct routing architectures for our study:
  • “Cos”: A gate that utilizes cosine similarity for routing, as detailed in (Chi et al. 2022). It is a two-layer architecture, with one linear layer projecting the features to a lower dimensional space and another performing cosine similarity against some learned latent codes.
We conducted training on the ImageNet-1K dataset using various small networks, each employing different routing mechanisms as displayed earlier. The results, as shown in Table 10, indicate that the simple convolutional (conv) configuration generally yields slightly better performance in most cases. However, when evaluating these models on out-of-distribution (OOD) datasets, the results are more varied. The inclusion of OOD evaluation is crucial as an effective routing mechanism should ideally facilitate an expert decomposition that generalizes well. . Notably, the MoE models that showed superior performance on ImageNet-1K also displayed enhanced robustness in the OOD evaluations.
ConvNext-S
ConvNext-S (iso.)
ViT-S
Appendix B Routing analysis
@@
Figure 3: The top row features the ConvNext-T-4 model, while the bottom row showcases the ViT-S-8 model, both of which have been trained on the ImageNet1k dataset. The images displayed have been generated by upsampling the gating output, retaining only the top expert for clarity. Each distinct colour represents a unique expert. These specific images are samples from the ImageNet validation set, extracted from the last third stage of the ConvNext-T-4 and the final layer of the ViT-S-8 models, respectively.

### v2

@@ 4.3.1 Impact of the Position of MoE Layers
Architecture MoE $\times 10^{6}$ Sample FLOPs IN-1K
@@
ı
Tab.˜3 displays the effects of various placements of MoE layers, comparing the following configurations: (a) Every 2: a MoE layer replaces every second layer; (b) Stage: a MoE layer replaces the final layer of each stage, resulting in four MoE layers throughout the network for ConvNeXt; (c) Last 2: a MoE layer replaces the final layer of each of the last two stages; (d) Last 3: Same as “Last 2”, but with an additional MoE layer inserted in the middle of stage 3. As shown in Tab.˜3, “Last 2” strategy is the most robust choice: it performs well across all architectures. In contrast, “Every 2” performs worst for ConvNeXt architecture, and the best for ViT. Consequently, in the following, all experiments use “Last 2” for ConvNeXt and “Every 2” for ViT.
MoE placement matters While ViT benefits most from inserting MoE layers at every second block and ConvNeXt from placing them in the last two stages, the Last 2 strategy offers a simple, broadly effective rule of thumb that consistently delivers strong performance across architectures.
4.3.2 Influence of the Number of Experts
@@
Architecture # experts Top-k #Params Per samples IN-1K
($\times 10^{6}$) #Paramsact Top 1 acc.
ConvNeXt-T no MoE - 28.6 28.6 82.1
4 1 54.3 25.5 82.1
8 1 47.4 25.5 82.1
8 2 47.4 25.5 82.0
ConvNeXt-T 16 1 46.3 25.5 81.7
ViT-S no MoE - 22.0 22.0 79.9
4 2 29.1 25.0 79.8
8 1 38.6 22.0 80.2
8 2 38.6 25.0 80.5
ViT-S 16 2 57.5 25.0 80.2
MoE #Params Per samples FLOPs IN-1K
strategy ($\times 10^{6}$) #Paramsact Top 1 acc.
@@
Tab.˜5 presents the performance of ConvNext and ViT architectures on ImageNet-1K training with the “Last 2” strategy MoE layers. It provides a comprehensive view of how the number of experts influences the size (parameter count) of the models and their Top-1 accuracy on this dataset.
The experimental results suggest that four experts yield the best performance for ConvNeXt, while eight experts are optimal for ViT. This is showcased by the Top-1 accuracy rates, which are consistent at 82.1% for ConvNeXt with four experts and reach a peak of 80.5% for ViT with eight experts. However, we note that increasing the number of experts to 16 has a detrimental effect on both architectures. Specifically, for ConvNeXt, the top-1 accuracy slightly drops to 81.7%, and for ViT, the performance plateaus at 80.2%.
These findings highlight that while MoE layers can enhance model performance, there is a delicate trade-off to be struck in terms of the number of experts. Exceeding the optimal number can lead to suboptimal results, negating the potential benefits of the MoE integration. Furthermore, in our detailed analysis, described in Sec.˜4.3.3, we explore the interplay between the number and the size of each expert. Our investigation revealed that, for isotropic networks such as ViT and ConvNeXt iso., reducing the size of experts adversely affects performance, while for ConvNeXt, it has no significant impact. Moreover, during the course of our experiments, we observed that while the Top-1 configuration was superior for ConvNeXt, the Top-2 configuration for ViT was comparable or even better than the ConvNeXt Top-1 configuration, leading us to employ Top-1 for ConvNeXt and Top-2 for ViT (cf. Tab.˜5).
Expert Count under Moderate Data Scale On ImageNet-1k, ConvNeXt models reach peak performance with 4 experts, while ViT benefits from up to 8. This contrast stems from architectural width: ConvNeXt’s broad final stages saturate quickly, limiting the value of additional experts, whereas ViT’s narrower last stage tolerates more experts.
4.3.3 Impact of Expert Size and Number
(a) # experts ($\in\{1,4,8,16\}$) vs. performance. Points’ size: MLP-ratio ($\in\{1,2,4\}$).
(b) Experts’ MLP-ratio ($\in\{2,4\}$) vs. performance. Points’ size: # experts ($\in\{1,4,8,16\}$).
Figure 3: Exploring the interplay between the size and count of experts in a MoE layer for ConvNeXt-T on ImageNet-1K. Baseline results (without MoE) are denoted by dotted lines. For this small dataset, MoE does not bring much improvement (see Fig.˜2 for bigger datasets).
Two critical tunable parameters in the MoE layers are the number of experts per layer and the size of each expert. The latter is defined by the mlp_ratio, which scales the hidden dimension of the feed-forward network (FFN) (Sec.˜3.1). Fig.˜3 illustrates the sensitivity of ConvNeXt’s performance to variations in these parameters, with the number of experts ranging from 4 to 16 and the mlp_ratio from 1 to 4. By default (original network), this ratio is set to 4, meaning that values below 4 correspond to smaller per-sample FFN capacities than in the original network.
To assess generalization, we report top-1 accuracy across three dataset categories: (i) in-distribution (ID) datasets—ImageNet and ImageNetV2 (Recht et al., 2019); (ii) out-of-distribution (OOD) datasets—ImageNet-R (Hendrycks et al., 2021a) and ImageNet-Sketch (Wang et al., 2019); and (iii) adversarial datasets—ImageNet-A (Hendrycks et al., 2021b). The objective is to determine whether performance benefits more from using many smaller experts or fewer larger ones.
Our experiments on ConvNeXt reveal that while MoE layers can slightly improve or preserve performance on ID datasets, they tend to degrade performance on OOD datasets. In particular, increasing the number of experts to 16 consistently reduces accuracy across all dataset types. Regarding expert size, smaller experts appear marginally more robust to increases in expert count. Overall, the best tradeoff for ID is achieved with 4 to 8 experts and an mlp_ratio of 2 or 4. Based on these findings, we adopt a configuration of 4 experts with an mlp_ratio of 2 for experiments on ImageNet-1K. For ConvNeXt isotropic and ViT architectures, however, reducing the mlp_ratio consistently harms performance, and in the case of ViT, it also introduces training instabilities. Also, OOD performance is typically degraded in Fig.˜3.
Balancing Expert Number and Width on ImageNet-1k On ImageNet-1k, the best trade-off is achieved with 4–8 experts and mlp_ratio 2–4; increasing to 16 experts consistently reduces accuracy and OOD robustness. Increasing number of experts while reducing their size helps reducing overfitting.
4.3.4 Results on ImageNet-21k
Tabs.˜5 and 6 present the results of models trained on ImageNet-21k, reporting ImageNet-1k accuracy after fine-tuning. We analyse the effect of MoE layer positioning, expert size (Tab.˜5), and expert count (Tab.˜6). The results demonstrate that, when large volumes of data are accessible, a greater number of experts can be effectively deployed. Notably, the use of sixteen experts, which was sub-optimal for ImageNet-1k, does not negatively affect performance on ImageNet-21k. This suggests that increasing the number of experts, together with the volume of data, could lead to valuable enhancements.
In terms of MoE layer configuration, our data showed no performance improvements with the addition of more layers, and the “every 2” strategy again yielded the poorest results for ConvNeXt-based architecture.
Scaling Expert Count with Larger Pretraining On ImageNet-21k, ConvNeXt scales up to 16 expert,s and ViT also benefits from more. Increased data reduces overfitting, allowing more experts. This highlights the link between dataset scale and MoE capacity. The optimal layer positioning remains unchanged.
Table 6: ImageNet-21k training on ConvNeXt-T: comparison between various numbers of experts, with an MLP ratio of 4.
@@ 4.3.4 Results on ImageNet-21k
4.3.5 Impact of the Routing Architectures
ConvNeXt-S
ConvNeXt-S (iso.)
ViT-S
We investigated three distinct routing architectures for our study:
  • “Cos”: A gate that utilizes cosine similarity for routing, as detailed in (Chi et al., 2022). It is a two-layer architecture, with one linear layer projecting the features to a lower-dimensional space and another performing cosine similarity against some learned latent codes.
We conducted training on the ImageNet-1K dataset using various small networks, each employing different routing mechanisms as displayed earlier. The results, as shown in Tab.˜7, indicate that the simple convolutional (conv) configuration generally yields slightly better performance in most cases. However, when evaluating these models on out-of-distribution (OOD) datasets, the results are more varied. The inclusion of OOD evaluation is crucial, as an effective routing mechanism should ideally facilitate an expert decomposition that generalizes well. Notably, the MoE models that showed superior performance on ImageNet-1K also displayed enhanced robustness in the OOD evaluations.
Simplicity Wins in Routing Architectures Across both ConvNeXt and ViT, a straightforward linear gate achieves the best overall results. More complex routing adds no clear benefit and often underperforms, even on OOD data.
5 Discussion
5.1 Hierarchical vs Isotropic Models
Table 8: Robustness of MoE w.r.t. to domain change, for two different pretrainings, both fine-tuned on ImageNet-1k: Top-1 accuracies on various datasets.
@@ 5.1 Hierarchical vs Isotropic Models
First, one can note that the best MoE designs typically apply experts only to the last layers. ViT operates at the same resolution at each layer after the input images have been cut into patches. This isotropic design implies that MoE is applied in layers that are not particularly large. On the other hand, with hierarchical models such as ConvNeXt, applying MoE to the last layers considerably increases model size. Indeed, as presented in Tab.˜6, employing eight experts doubles the size of the network despite using only two layers of MoE.
5.2 Positions and Numbers of MoE Layers
There are two prevalent strategies in the literature concerning the position of MoE layer: “Every 2” and “Last 2”. In the context of hierarchical architectures such as ConvNeXt, the “Every 2” strategy dramatically increases the number of weights while simultaneously yielding inferior results. However, for ViT “Last 2” and “Every 2” are two viable strategies. This is in line with the results of V-MoE (Riquelme et al., 2021) on large-scale datasets. Overall, our results confirm that “Last 2” is a solid starting point, as it consistently yields positive outcomes across all tested architectures. However, it may not always be the optimal choice for every architecture. For instance, in the case of ViT, “Every 2” demonstrates better results. Regarding the number of experts, we noted that model performance quickly saturates with the number of experts, as reproduced in Tab.˜6, and detailed in Fig.˜3.
5.3 Robustness Evaluation
We evaluated MoE models, trained or fine-tuned on ImageNet-1k, against a range of robustness benchmark datasets, such as ImageNet-A (IMA) (Hendrycks et al., 2021b), ImageNet-R (IMR) (Hendrycks et al., 2021a), ImageNet-Sketch (IMSK) (Wang et al., 2019), and ImageNet V2 (IMV2) (Recht et al., 2019), the latter being used as a measure of overfitting.
@@ 5.3 Robustness Evaluation
Tab.˜8 reports Top-1 accuracy for all datasets. It shows that MoE models tend to underperform their dense counterparts in ImageNet-1k. However, we note that MoE models are better for the robustness metric; they were actually already better on ImageNet-1k. So, these results actually emphasize the success of MoE when the model is sufficiently small to be improved by a MoE, as discussed in the previous section: the domain generalization holds only in the regime in which MoE improves the base results.
5.4 On Data-Augmentation
As shown in the previous section, MoE is sensitive to the amount of data at hand. This sensitivity stems from the inherent structure of MoE, where increasing the number of experts results in each expert operating on a smaller data fraction. Consequently, each expert tends to learn specific data clusters, as discussed by (Chen et al., 2022).
Modern training methodologies often employ robust data augmentation techniques to enhance model generalization. However, the interplay between these augmentations and the data clusters learned by the MoE model remains relatively uncharted. For instance, techniques like Mixup (Zhang et al., 2018), which blend images, could blur the distinctions between these clusters. Similarly, the Random Resize method, with its aggressive magnification capability, may distort the original data clusters. Understanding the precise influence of these augmentations on the MoE model’s learning process is essential, as it holds implications for the model’s ultimate performance. Some evidence of this can be seen in Tab.˜9. While one training recipe might show MoE outperforming its dense counterparts (as indicated in the top rows of the table), a different, more effective recipe could lead to the opposite outcome (Bottom rows of the table).
Table 9: ViT-B trained on ImageNet-1k with different hyperparameters. In particular, for the top row (gray) we follow the training recipe from Deit (Touvron et al., 2021) and for the bottom one we follow Deit III (Touvron et al., 2022).
Architecture # Params Per sample IN-1K
($\times 10^{6}$) #Paramsact Acc.
ViT-B-8
(Touvron et al., 2021) 86.6 86.6 81.8
ViT-B-8
Every 2 Top 2 284.9 129.9 82.2
ViT-B
(Touvron et al., 2022) 86.6 86.6 82.8
ViT-B-8
Every 2 Top 2 284.9 129.9 82.5
Figure 4: Improvement from MoE (y-axis) vs. baseline accuracy without MoE (x-axis) for image classification, across pretraining sizes. Larger pretraining yields greater gains; higher baseline accuracy reduces impact. Dashed lines show maximal improvement per accuracy level. Results are drawn from Tabs.˜2, 12 and 11 (in App.˜B).
5.5 Alternate Views
With the ImageNet benchmark increasingly reaching saturation, the focus in contemporary computer vision models has shifted towards computational efficiency and scaling prowess. An essential question arises: Does the MoE model enhance this aspect in image classification? Our experiments with ImageNet reveal an intriguing insight: integrating MoE into models like ConvNeXt and ViT enhances performance, but this improvement tends to plateau in models with over 100M parameters. At first glance, this might seem to contradict findings from some papers (Riquelme et al., 2021; Mustafa et al., 2022). However, with additional context, this observation aligns with the existing literature, as evidenced by results in Figs.˜4 and 2, and Tab.˜11 (App.˜B). Specifically, as the base dense network grows larger, the performance gap between MoE and its dense counterpart narrows.
Exploring this further in large-scale datasets, such as JFT 3 billion (Zhai et al., 2022), reveals similar findings. For instance, the vision transformer model SoViT, with 400M parameters (Alabdulmohsin et al., 2023), exhibits comparable performance to the 15B V-MoE (Riquelme et al., 2021) when both use ViT-based architectures. Similarly, when considering the JFT-3B dataset (Zhai et al., 2022), the 5.6B LiMoE-H model (Mustafa et al., 2022) and Coca-Large (Yu et al., 2022) lead to analogous results when evaluated on a per-sample parameter basis. Interestingly, in the vision-language modeling domain, Lin et al. (2024) observe that increasing the number of experts for the image modality does not lead to diminishing returns, aligning with our findings regarding the efficacy of experts in scaling performance.
To summarize, while the MoE model does not appear to redefine the state-of-the-art when matched against dense models on both ImageNet and billion-scale datasets, its real strength emerges in enhancing smaller models. By integrating MoE, these compact models can achieve better performance, pushing their capabilities closer to the forefront of current benchmarks. Yet, it is essential to acknowledge that, in the broader context, MoE might not drive significant advancements in the overall state-of-the-art. Moreover, shape optimization as demonstrated in SoViT, could further improve performance and could even compound with the benefits of MoE, enabling smaller and more capable models. This nuanced perspective is further enforced in the appendix, with the data showcased in Tab.˜11 in App.˜B.
5.6 Model Inspection
Figure 5: Distribution of experts
@@ 5.6 Model Inspection
The objective of this section is to examine the routing component of the model. We will present the top expert for each routing layer. Note that when considering the top-2 configuration, we only report, for this visualization, the top-1 of these two experts.
@@
Figure 6: The top row features the ConvNext-T-4 model, while the bottom row showcases the ViT-S-8 model, both of which have been trained on the ImageNet1k dataset. The images displayed have been generated by upsampling the gating output, retaining only the top expert for clarity. Each distinct color represents a unique expert. These specific images are samples from the ImageNet validation set, extracted from the last third stage of the ConvNeXt-T-4 and the final layer of the ViT-S-8 models, respectively.
(a) ViT-B deepest MoE. No absolute positional bias is observed.
(b) ConvNeXt-B last MoE layer. A clear spatial partitioning of experts emerges.
Figure 7: Average spatial distribution of experts trained on ImageNet-22K.
Visual inspection: Fig.˜6 provides a visualization of the portions of images that are assigned to the different experts. This reveals that some experts specialize in specific elements of an image, such as animals, buildings, or objects. This trend is especially apparent when considering the top-1 MoE, but the distinction becomes less clear with the top-2 MoE. In both instances, the clusters formed by each expert are challenging to interpret, and the spatial locations of each expert tend to lack clarity and can be quite ambiguous. However, when directly examining the spatial distribution, a clear spatial decomposition emerges for ConvNeXt models (cf. Fig.˜7). In particular, Fig.˜7(b) shows that certain experts focus specifically on the image borders or the center. This behavior may be explained by the strong local spatial bias imposed by convolution, in contrast to the attention-based models, which do not exhibit such patterns.
Number of experts involved per image: We quantified the average number of experts contributing to processing an image. In the initial layers, the majority of experts are typically engaged for each image. With fewer than eight experts, most are utilized, even in deeper layers. For instance, in the ViT model, with eight experts, an average of seven experts contributed per image, while in the ConvNeXt model, there are six. As we increase the number of available experts per MoE layer, the number of contributing experts per image also rises. For example, in ConvNeXt, with 16 experts, an average of 10.5 experts are involved, and with 32 experts, this number increases to 15.8. This suggests that each expert is likely focused on a limited number of image patches. This observation is confirmed by the cumulative distribution function (CDF) of each expert per image, as depicted in Fig.˜8. On the other hand, when directly examining the routing distribution across experts, we observe a well-balanced assignment, indicating that the load-balancing loss is effectively fulfilling its role, as depicted in Fig.˜5.
(a) 8 experts
(b) 16 experts
(c) 32 experts
Figure 8: Cumulated Distribution Function of the number of patches used per image for ConvNeXt-S trained on ImageNet-21k, with 8, 16, and 32 experts. Each colour corresponds to one expert. Sample detailed interpretation: for 8 experts, the “red” expert sees between 1 and 30 patches (out of 196) for 80% of the images on which it is active. In particular (and even more so when there are many experts), an expert frequently covers a limited part of the image. For instance, if the y-axis is close to 1 at x-axis=100 patches, covering more than 100 patches over 196 is rare: experts don’t specialize much. As the number of experts increases, this tendency of experts being used only on a few patches increases too. For 32 experts, most of them focus on less than 30 patches; 80% of the time, they are active.
Correlation between experts and labels: Fig.˜9 (top) presents the correlation between the number of expert occurrences and individual classes in ImageNet-1k. The class IDs in ImageNet-1k are organized in such a way that adjacent classes often share similar attributes. For instance, class IDs ranging from 151 to 268 are all dedicated to different dog breeds. From Fig.˜9, we note that the occurrences in the first three MoE layers do not match the ImageNet-1k classes, with most experts appearing uniformly across different classes. However, in the last three MoE layers, there is a noticeable trend of more experts aligning with specific ImageNet-1k classes. This distinction is especially pronounced between experts focusing on animals (classes up to ID 397, such as experts 0 and 6 in Fig.˜9-(e)) and those centered on objects (classes beyond ID 397, like experts 1 and 5 in Fig.˜9-(e)). As the number of experts increases, this class-specific alignment still holds, though it remains challenging to clearly define the roles of these experts since they often span several scattered classes rather than forming tight, contiguous groups that would fit each class, as illustrated in Fig.˜10.